# Artificial Human Intelligence: The role of Humans in the Development of Next Generation AI

### v1

@@ 3.2.1 Human–inspired AI: Neuroscience meets Machines
Machines can be inspired by humans in various conspicuous ways. For instance, humans are recently shown to perform blur-detection tasks on faces using two different dominant strategies [45]. These strategies are under the influence of various factors including but not limited to top-down processes, bias, and behavioral similarities or differences. Behavioral characterizations of such human proclivities [46, 47] can be quantified and transferred to machine implementations to advance their guessing strength for multiple-choice tasks. Guessing framework is mathematically sound and inspiring piece of tool to design as well as evaluate neural network strategies with the potential of optimal guessing techniques being incorporated as part of learning [48, 49]. Moreover, psychophysics and neural activations can be used to understand the temporal ordering of critical processes and spatial distribution of basic functions, such as perception, performed by specialized cortical circuits. The results of said studies are essential to choosing the right inductive biases, architecture as well as learning (optimization) schemes for embedding human decision-making processes as part of the model development. It is quite a fortune that the human brain can region-specifically be monitored while actively functioning through information extraction techniques based on electrical or magnetic principles. The neuroimaging and electrophysiological methodologies that come in handy in the analysis of the human brain include EEG [50], EMG, MEG [51], fMRI [52], fNIRS [53], etc. Overall, these techniques can enhance our understanding of the modular and specialized nature of the brain, causal order of temporal and spatial processing, and ubiquitous muscle functions, facilitating comprehensive analyses that advance neuroscience, psychology, medicine, and related fields. However, there is no known commensurate tool to measure similar features in deep networks (e.g. modularity) despite a few solid attempts that have been made in the information and statistics theory domain [54, 55]. The main objective of these mathematical tools is to primarily measure information modularity and evaluate robustness which results in leveraging algorithmic techniques to force models to demonstrate similar characteristics concerning processing specializations.
@@
One of the most fascinating phases of human brain development is postnatal period, covering early few years of infancy where new brain connectivity suddenly transpires at insane rates in light of neural prolifiration and subsequent prunning processes [56]. This dynamic mission of architectural modifications can be shown to be beneficial even in normal brain development [57]. In the meantime, frontend optics as well as retinal anatomy also develop and start allowing more light to enter and higher number of cells to participate in retinal image processing, respectively, for more precise spiking, higher-level and accurate perception. It is shown in several past studies that poor visual acuity in human newborns offers several advantages, including faster and more accurate acquisition of visual object categories at multiple hierarchical levels [58]. These structural changes in every part of visual system such as chromatic and contrast sensitivities and its impact on the backend cortical substrate and functional connectivity can be used to design biologically-plausible (or so called biomimetic) training regimens to help train the machines in a similar way [58, 59, 60]. For instance, a training regimen that begins with high blur and gradually reduces its effect over time (allowing more high frequency to gradually enter), mimicking human development, is shown to enhance basic-level categorization performance compared to using non-blurred inputs throughout the training. These progressions may hold adaptive value by fostering the formation of receptive field structures that confer resilience to subsequent spatial or chromatic degradations. Fig. 4 demonstrates an example of biomimetic progression of acuity and color, applied to a toy image. As can be seen while the initial days of post natal period sees black and white, the maturation of cones leads to gradual processing of color later in development. However, in human development, not only are spatial frequencies and color subject to dynamic changes, but also the anatomical structure, generated internal noise over time, and maturation of recurrent pathways also contribute to this biomimetic progression. These factors collectively perfect the input feedback-consolidated architectural refinements over time, more so in the infancy [61, 62, 63].
@@
One of the well-known barriers to good generalization is the overfitting phenomenon in machines. To practically combat against overfitting, regularization is proposed by injecting noise into the system to ensure bare minimum robustness [64, 65] or expedite the training time [66]. The addition of noise into training is quite similar in spirit to noisy optimizations within the context of simulated annealing [67, 68], which can help the model escape local minima by allowing it to explore a broader region in the loss landscape. By introducing noise, the optimization process can avoid getting trapped in small basins and suboptimal points, thereby improving the chances of finding a potential global minimum located inside a large basin. This mechanism is particularly useful in deep neural networks, where the loss landscape is highly non-convex with many local minima [69]. Understanding the learning mechanisms in humans, particularly the effects of attention and focus on the learning process, is crucial. These biological phenomena parallel computational techniques that mitigate erroneous learning and facilitate fact-checking, thereby ensuring the accuracy of information processed by the belief subsystem.
@@
3.2.2 Human–assisted AI: Hybrid Systems
@@
In Human-assisted AI, models are developed with the primary goal of augmenting or enhancing human/machine capabilities rather than replicating machine or human intelligence. Human-assisted AI leverages AI technologies to assist machines or humans in performing tasks more efficiently, accurately, or autonomously. Human–assisted AI, as a bidirectional enterprise, can be characterized as the continuous interaction between human-helping-machines and machines-helping-humans. However, fully realizing the potential of human-AI collaboration requires to address important challenges. First, the necessary conditions that support complementarity – where human-assisted AI performance is expected to surpass that of AI alone for a given task, requiring humans to recognize when to augment AI that complement critical decision-making processes. Second, accurately assessing human mental models as well as the extent to which humans are prone to failures are crucial, which also brings out the issue of understanding the effects of design choices in human-AI interaction, such as the timing and amount of Human assistance to avoid bias.
@@
Primary example of this trend is the foundation models that can interact with humans, such as most recent advanced LLMs [74] (GPT4 [6], Llama [75], etc.) that begin transforming the software industry by providing assistance writing the bulk of the required code as a downstream task, and still subject to thorough validation [76, 77]. This approach (especially the former reference [76]) emphasizes the significance of collaboration between humans and machines, with AI systems providing support, guidance, or automation to complement human skills (critical thinking, reasoning, navigation of complex decision processes) and expertise. Other examples include AI-augmented tools for object recognition, memory enhancement [78], data analysis, medical diagnosis [79], language translation, and robotic assistance in manufacturing or logistics [80]. Yet another hot topic could be mission critical applications where humans and machines could perform tasks using different strategies and hence end up having distinct failure patterns. This complementarity can be utilized as a fusion to bolster overall accuracy performance [81].
@@
Human mind organization and sensorimotor systems could be useful for developing computational systems in various ways. One remarkable example from vision is the leverage of eye tracking and analysis of gaze direction, which has the potential to uncover a wealth of information about internal cognitive states. For example, it is argued that gaze can reveal the underlying attentional patterns [82]. More broadly, gaze information could be useful from two standpoints. It can help AI develop an attention mechanism to effectively cope with the information-rich world and determine significant cognitive features [83]. In fact, given the amount of training data required to learn such an adaptive attention process, it may be easier for AI agents to learn attention directly from human gaze data. Secondly, it can help AI interfaces to perceive and understand gaze movements to create naturalistic interactions with humans [84, 85]. For instance, gaze-based control mechanisms could allow users to interact with AI systems without physical input devices (keyboards, mouse, etc.), fostering a more immersive and responsive interaction.
@@
3.2.3 Human–independent AI: The rise of Open Intelligence
@@
In our taxonomy, this branch is still considered as human-centered AI because the studies in this category are still framed based on human intelligence as a reference model. For instance, although AlexNet [34] is inspired by the visual cortex and its layered functional structure, designed to mimic human visual perception and processing, the proposed optimization algorithm backpropagation (also referred as back-prop) is not necessarily human-inspired; it is more inspired by how computational constraints shape and efficiently utilize resources, along with mathematical convenience such as sticking to the same set of weights/coefficients in both forward and backward passes [86]. In fact, back-prop is considered to be an optimization method which spreads information across all parameters of the neural network making the network less modular [55] (also see Section 4.3). Hence one can argue that the development of AlexNet is not necessarily categorized as human-inspired but rather neuroscience-inspired at the very high-levels of abstraction subject to SoTA silicon hardware constraints and explainability.
@@
The main motivation behind Human-independent AI is twofold: (1) In pursuit of an open intelligence that is independent of human intelligence could potentially be a better model for the natural world we live in as human intelligence seems to have missing elements to tackle important issues of the world such as less advanced predictive capabilities for extended time and range or the inherent limitations of convolution/pooling operations executed by the visual system. This is corroborated by the most recent transformer models which lifts off some of the important inductive biases that previous generation CNNs adopted. Second observation to this idea is the ever growing data scales as machines having less inductive bias would require to build such biases based on the voluminous data. (2) In any reliable/efficient AI system development, we would expect a joint symbiotic life between the intelligence and the hardware that runs it. The current silicon hardware and Von-Neuman architecture is not necessarily the right home for the optimal implementation of human intelligence. This might guide us to think of other forms of intelligences that could optimally run on the available hardware architectures, building on the few-decade experience of distributed microprocessor history of human construct. For more elaborate discussion on the effect of hardware on the implemented intelligence, please see 5.3.1.
@@
Perhaps it might be more accurate to frame human-independent AI as an effort which characterizes humans as irrelevant to the next generation large-scale AI development. There is a growing inclination in the community believing that the independent design will benefit from some of the findings that come out through interactions with the outside world (experimentations etc.) which might well be counter-intuitive and cannot be adequately represented by the data intrinsics. However, these designs are still bound to subjective evaluations by human-credited metrics and might be compared to human performance to improve their explainability. After all, explainability is a human-construct and only meaningful if it helps with understanding the true nature of phenomenon.
@@
4 Brain-inspired Information processing
@@
Brain-Inspired Information Processing refers to building computational systems and designing architectures that emulate the brain’s structure and functions to achieve advanced cognitive and data processing abilities. This interdisciplinary field draws on neuroscience, cognitive science, and artificial intelligence to develop models and systems that implement brain-like processes for tasks like multi-modal perception, learning in development, and critical decision-making/building commonsense. The overall picture is illustrated in Fig. 5 where the elements of brain-inspired information processing as well as the context is introduced in which this paradigm is utilized [87]. To get a grasp on how the brain operates, various neuroimaging techniques are utilized, including but not limited to Electroencephalography (EEG) and Magnetoencephalography (MEG) which measure the brain’s electrical and magnetic activity, offering high temporal resolution for tracking dynamic neural activity [88]. Functional Magnetic Resonance Imaging (fMRI), on the other hand, provides detailed spatial resolution by monitoring blood flow changes, helping to map which brain regions are involved in particular cognitive functions [89]. These tools are crucial for revealing the neural dynamics and connections that underpin complex cognitive processes. To identify neural signatures for different brain functions, machine learning algorithms are trained on brain signal patterns associated with specific tasks or mental states. By analyzing the relationships between brain regions, statistical methods (such as random graph theory) help model brain networks and pinpoint crucial regions/nodes involved in processing information [90].
@@
The informative features extracted from the biomedical data are typically used to build computational models that mimic the way our brains process information. Translating these measurements into computational frameworks that are truly brain-inspired could be quite challenging though. For example, neural networks can be structured to reflect the hierarchical nature of the brain, with different layers corresponding to various levels of processing found in, say, the visual cortex [91, 92, 89]. Finally, these computational frameworks are refined by comparing them against real brain data, perhaps at higher abstraction levels, ensuring they not only mimic brain activity but also can generalize to new, unseen tasks. However, the resemblance is always subject to questions as it is hard to completely translate brain operations to algorithms and implement them on silicon. By blending neuroimaging with AI and statistical analysis, researchers mainly aim to create brain-inspired models that capture, at least to some extent, the core manifestation of human cognition. These models can then be applied to fields ranging from the most popular foundation models such as LLMs to autonomous systems. In fact, brain-inspired data processing could guide foundation models to embrace emergent behaviours more rapidly as architectural similarity is more likely to demonstrate functional similarities [93] (See also Section 7 for further discussions). This approach seeks to bridge the gap between biological intelligence and machine intelligence, pushing toward more flexible and robust AI systems.
To fully adapt brain-inspired processing, it is crucial to tackle several fundamental questions, which can be grouped, to our count, into four key areas: (1) the possibility of reverse engineering of human cognitive abilities, (2) the distinctive features of the human brain, (3) the mechanisms through which our brains achieve modularity and robustness, and (4) exploration of gaps in current state-of-the-art approaches that prevent us from reaching human-level AI capabilities.
4.1 Reverse engineering a human skill

### v2

@@ III-C Comparisons between HIns, HAss and HInd Intelligences
III-C1 HIns Intelligence (Neuroscience meets Machines)
In Human-inspired AI, machines are designed to mimic or replicate aspects of human intelligence, behavior, or complex cognitive processes. Part of this endevour could be inspired by the brain mechanisms responsible for processing information and navigating attention to complete complex tasks. Human-inspired AI aims at understanding and replicating the neural mechanisms behind human cognitive processes, such as learning, common sense reasoning, perception, and critical decision-making. This approach often involves studying biological systems closely including neuroscience, psychology, cellular chemistry and cognitive sciences to accurately inform and guide the design of AI algorithms and architectures. Perhaps the most critical phase of the formulation of human–like intelligence is the developmental phases of the brain anatomy and its training (learning trajectory) with psychophysical limitations impacting its evolution. Examples include neural networks inspired by the brain’s architecture (structural and functional connectivity) and specialization of the specific human brain regions, biologically plausible learning algorithms [44] or algorithms (recursive, top-down and bottom-up) that accurately model human decision-making processes. Human-inspired AI can be designed through methodological changes in model architecture or training. Architectural changes could be implemented through structural or functional changes (connectivity and topology) involving addition or removal of nuronal units in the process of training/learning processes.
@@ III-C1 HIns Intelligence (Neuroscience meets Machines)
Machines can be inspired by humans in various conspicuous ways. For instance, humans are recently shown to perform blur-detection tasks on faces using two different dominant strategies [45]. These strategies are under the influence of various factors including but not limited to top-down processes, bias, and behavioral similarities or differences. Behavioral characterizations of such human proclivities [46, 47] can be quantified and transferred to machine implementations to advance their guessing strength for multiple-choice tasks. Guessing framework is mathematically sound and inspiring piece of tool to design as well as evaluate neural network strategies with the potential of optimal guessing techniques being incorporated as part of learning [48, 49]. Moreover, psychophysics and neural activations can be used to understand the temporal ordering of critical processes and spatial distribution of basic functions, such as perception, performed by specialized cortical circuits. The results of said studies are essential to choosing the right inductive biases, architecture as well as learning (optimization) schemes for embedding human decision-making processes as part of the model development (model bias). It is quite a fortune that the human brain can region-specifically be monitored while actively functioning through information extraction techniques based on electrical or magnetic principles. The neuroimaging and electrophysiological methodologies that come in handy in the analysis of the human brain include EEG [50], EMG, MEG [51], fMRI [52], fNIRS [53], etc. Overall, these techniques can enhance our understanding of the modular and specialized nature of the brain, causal order of temporal and spatial processing, and ubiquitous muscle functions, facilitating comprehensive analyses that advance neuroscience, psychology, medicine, and related fields. However, there is no known commensurate tool to measure similar features in deep networks (e.g. modularity) despite a few solid attempts that have been made in the information and statistics theory domain [54, 55]. The main objective of these mathematical tools is to primarily measure information modularity and evaluate robustness which results in leveraging algorithmic techniques to force models to demonstrate similar characteristics concerning processing specializations.
@@
One of the most fascinating phases of human brain development is postnatal period, covering early few years of infancy where new brain connectivity suddenly transpires at insane rates in light of neural prolifiration and subsequent prunning processes [56]. This dynamic mission of architectural modifications can be shown to be beneficial even in normal brain development [57]. In the meantime, frontend optics as well as retinal anatomy also develop and start allowing more light to enter and higher number of cells to participate in retinal image processing, respectively, for more precise spiking, higher-level and accurate perception. It is shown in several past studies that poor visual acuity in human newborns offers several advantages, including faster and more accurate acquisition of visual object categories at multiple hierarchical levels [58]. These structural changes in every part of visual system such as chromatic and contrast sensitivities and its impact on the backend cortical substrate and functional connectivity can be used to design biologically-plausible (or so called biomimetic) training regimens to help train the machines in a similar way [58, 59, 60]. For instance, a training regimen that begins with high blur (low spatial resolution) and gradually reducing its effect over time (allowing more high spatial frequency to gradually enter), mimicking human vision development, is shown to enhance basic-level categorization performance compared to using full resolution image inputs throughout the training. These progressions may hold adaptive value by fostering the formation of receptive field structures that confer resilience to subsequent spatial or chromatic degradations. Fig. 5 demonstrates an example of biomimetic progression of acuity and color, applied to a toy image. As can be seen while the initial days of post natal period sees black and white, the maturation of cones leads to gradual processing of colors later in development. However, in vision development, not only are spatial frequencies and color subject to dynamic changes, but also the anatomical structure, generated internal noise over time, and maturation of recurrent pathways also contribute to this biomimetic progression. These factors collectively perfect the input feedback-consolidated architectural refinements over time, more so in the infancy [61, 62, 63].
@@
One of the well-known barriers to good generalization is the overfitting phenomenon in machines. To practically combat against overfitting, regularization is proposed by injecting noise into the system to ensure bare minimum robustness [64, 65] or expedite the training time [66]. The addition of noise into training is quite similar in spirit to noisy optimizations within the context of simulated annealing [67, 68], which can help the model escape local minima by allowing it to explore a broader region in the loss landscape. By introducing noise (or some form of stochasticity), the optimization process can avoid getting trapped in small basins and suboptimal points, thereby improving the chances of finding a potential global minimum located inside a large basin. This mechanism is particularly useful in deep neural networks, where the loss landscape is highly non-convex with many local minima [69]. Understanding the learning mechanisms in humans, particularly the effects of attention and focus on the learning process, is crucial. These biological phenomena parallel computational techniques that mitigate erroneous learning and facilitate fact-checking, thereby ensuring the accuracy of information processed by the belief subsystem.
@@
III-C2 HAss Intelligence (Birth of Hybrid Systems)
@@
In Human-assisted AI, models are developed with the primary goal of augmenting or enhancing human/machine capabilities rather than replicating machine or human intelligence. Human-assisted AI leverages AI technologies to assist machines or humans in performing tasks more efficiently, accurately, or autonomously. Human–assisted AI, as a bidirectional enterprise, can be characterized as the continuous interaction between human-helping-machines and machines-helping-humans. However, fully realizing the potential of Human–AI collaboration requires to address important challenges. First, the necessary conditions that support complementarity – where human–assisted AI performance is expected to surpass that of AI alone for a given task, requiring humans to recognize when to augment AI that complement critical decision–making processes. Second, accurately assessing human mental models as well as the extent to which humans are prone to failures are crucial, which also brings out the issue of understanding the effects of design choices in human–AI interaction, such as the timing and amount of Human assistance to avoid bias.
@@
Primary example of this trend is the foundation models that can interact with humans, such as most recent advanced LLMs [74] (GPT4 [6], Llama [75], etc.) that begin transforming the software industry by providing assistance writing the bulk of the required source code as a downstream task, and still subject to thorough validation [76, 77]. This approach (especially the former reference [76]) emphasizes the significance of collaboration between humans and machines, with AI systems providing support, guidance, or automation to complement human skills (critical thinking, reasoning, navigation of complex decision processes) and expertise. Other examples include AI–augmented tools for object recognition, memory enhancement [78], data analysis, medical diagnosis [79], language translation, and robotic assistance in manufacturing or logistics [80]. Yet another hot topic could be mission critical applications where humans and machines could perform tasks using different strategies and hence end up having distinct failure patterns. This complementarity can be utilized as a fusion to bolster overall accuracy performance [81].
@@
Human mind organization and sensorimotor systems could be useful for developing computational systems in various ways. One remarkable example from Human vision is the leverage of eye tracking and analysis of gaze direction, which has the potential to uncover a wealth of information about internal cognitive states. For example, it is argued that gaze can reveal the underlying attentional patterns [82]. More broadly, gaze information could be useful from two standpoints. It can help AI develop an attention mechanism to effectively cope with the information-rich world and determine significant cognitive features [83]. In fact, given the amount of training data required to learn such an adaptive attention process, it may be easier for AI agents to learn attention directly from human gaze data. Secondly, it can help AI interfaces to perceive and understand gaze movements to create naturalistic interactions with humans [84, 85]. For instance, gaze-based control mechanisms could allow users to interact with AI systems without physical input devices (keyboards, mouse, etc.), fostering a more immersive and responsive interaction.
@@
III-C3 HInd Intelligence (Algorithms for Machine Hardware)
@@
In our taxonomy, this branch is still considered as human-centered AI because the studies in this category are still framed based on human intelligence as a reference model. For instance, although AlexNet [34] is inspired by the visual cortex and its layered functional structure, designed to mimic human visual perception and processing, the proposed optimization algorithm backpropagation (also referred as back-prop) is not necessarily human-inspired; it is more inspired by how computational constraints shape and efficiently utilize resources, along with mathematical convenience such as sticking to the same set of weights/coefficients in both forward and backward passes [86]. In fact, back-prop is considered to be an optimization method which spreads information across all parameters of the neural network making the network less modular [55]. Hence one can argue that the development of AlexNet is not necessarily categorized as human-inspired but rather neuroscience-inspired at the very high-levels of abstraction subject to SoTA silicon hardware constraints and explainability.
@@
The main motivation behind Human-Independent AI is twofold: (1) In pursuit of an open intelligence that is independent of human intelligence could potentially be a better model for the natural world we live in as human intelligence seems to have missing elements to tackle important issues of the world such as less advanced predictive capabilities for extended time and range or the inherent limitations of convolution/pooling operations executed by the visual system. This is corroborated by the most recent transformer models which lifts off some of the important inductive biases that previous generation CNNs adopted. Second observation to this idea is the ever growing data scales as machines having less inductive bias would require to build such biases automatically based on the peculiarities of voluminous data. (2) In any reliable/efficient AI system development, we would expect a joint symbiotic life between the intelligence and the hardware that runs it. The current silicon hardware and Von-Neuman architecture is not necessarily the right home for the optimal implementation of human intelligence. This might guide us to think of other forms of intelligences that could optimally run on the available hardware architectures, building on the few-decade experience of distributed microprocessor history of human construct. For more elaborate discussion on the effect of hardware on the implemented intelligence, please see IV-D1.
@@
Perhaps it might be more accurate to frame human-independent AI as an effort which characterizes humans as irrelevant to the next generation large–scale AI development. There is a growing inclination in the community believing that the independent design will benefit from some of the findings that come out through interactions with the outside world (experiments etc.) which might well be counter-intuitive and cannot be adequately represented by the data intrinsics. However, these designs are still bound to subjective evaluations by human–credited metrics and might be compared to human performance to improve their explainability. After all, explainability is a human-construct and only meaningful if it helps with understanding the true nature of phenomenon.
@@
IV Human-Level AI and Challenges/Perspectives
By combining insights from human cognition (human-inspired) and biological neural structures (brain-inspired), human-level AI strives to achieve general intelligence while maintaining ethical, contextual, and adaptive capabilities similar to humans.
IV-A Brain-inspired Information processing
@@
Brain-Inspired Information Processing, ranked at the top of intelligence inspiration tree presented in Fig. 1, refers to building computational systems and designing architectures that emulate the brain’s structure and functions to achieve advanced cognitive and data processing abilities. This interdisciplinary field draws on neuroscience, cognitive science, and artificial intelligence to develop models and systems that implement brain-like processes for tasks like multi-modal perception, learning in development, and critical decision-making/building commonsense. The overall picture is illustrated in Fig. 6 where the elements of brain-inspired information processing as well as the context is introduced in which this paradigm is utilized [87]. To get a grasp on how the brain operates, various neuroimaging techniques are utilized, including but not limited to Electroencephalography (EEG) and Magnetoencephalography (MEG) which measure the brain’s electrical and magnetic activity, offering high temporal resolution for tracking dynamic neural activity [88]. Functional Magnetic Resonance Imaging (fMRI), on the other hand, provides detailed spatial resolution by monitoring blood flow changes, helping to map which brain regions are involved in particular cognitive functions [89]. These tools are crucial for revealing the neural dynamics and connections that underpin complex cognitive processes. To identify neural signatures for different brain functions, machine learning algorithms are trained on brain signal patterns associated with specific tasks or mental states. By analyzing the relationships between brain regions, statistical methods (such as random graph theory) help model brain networks and pinpoint crucial regions/nodes involved in processing information [90].
@@
The informative features extracted from the biomedical data are typically used to build computational models that mimic the way our brains process information. Translating these measurements into computational frameworks that are truly brain-inspired could be quite challenging though. For example, neural networks can be structured to reflect the hierarchical nature of the brain, with different layers corresponding to various levels of processing found in, say, the visual cortex [91, 92, 89]. Finally, these computational frameworks are refined by comparing them against real brain data, perhaps at higher abstraction levels, ensuring they not only mimic brain activity but also can generalize to new, unseen tasks. However, the resemblance is always subject to questions as it is hard to completely translate brain operations to algorithms and implement them on silicon. By blending neuroimaging with AI and statistical analysis, researchers mainly aim to create brain-inspired models that capture, at least to some extent, the core manifestation of human-level cognition. These models can then be applied to fields ranging from the most popular foundation models such as LLMs to autonomous systems. In fact, brain-inspired data processing could guide foundation models to embrace emergent behaviours more rapidly as architectural similarity is more likely to demonstrate functional similarities [93] (See also Section VI for further discussions).
IV-A1 Reverse engineering a human skill